# Material Point Methods on Unstructured Tessellations: A Stable Kernel Approach With Continuous Gradient Reconstruction

### v1

@@ 2.3.2 Extending MLS-MPM Onto Unstructured Meshes
We select MLS-MPM as the foundational approach due to the inherent versatility of MLS, which enables application to adjacent nodes without reliance on specific topological or positional constraints. Our implementation and experimental work have focused on triangular and tetrahedral meshes. Nonetheless, it is worth noting that our method can easily be extended to other kinds of unstructured grid tessellations by designing a smooth and locally diminishing function $\eta_{v}$ compatible with the tessellation, such as the one we design below in (23) for simplex cells.
@@
To determine the nearby vertices for a given particle $p$, we first locate the cell that encompasses $p$ and refer to its vertices as ${\mathcal{N}}_{p}^{0}$, representing the 0-ring neighbors of $p$. Then, we define ${\mathcal{N}}_{p}^{1}$ as the 1-ring neighbors, which comprise all nodes connected to ${\mathcal{N}}_{p}^{0}$. Note that ${\mathcal{N}}_{p}^{0}$ is a subset of ${\mathcal{N}}_{p}^{1}$. Similarly, we can define ${\mathcal{N}}_{p}^{2},\dots$ in an analogous manner, as illustrated in Figure 1.
Figure 2: (a) When ${\mathcal{N}}_{p}^{0}$ alone is selected as the active Degrees of Freedom (DOFs), as a particle crosses the cell edge, the DOFs indicated by the blue and red boxes are added or removed, respectively. Consequently, the weights on these DOFs must approach zero to ensure ${\mathcal{C}}^{0}$ continuity, resulting in kernel degeneration along the edge. (b) Advancing to ${\mathcal{C}}^{1}$ addresses this issue by incorporating a sufficient number of DOFs to fully encompass the particles.
@@
When a specific level of ring neighbors is chosen as the nearby nodal degrees of freedom, a natural question arises: What is the minimum number of rings required to satisfy the desired properties of the MPM kernel? Assume ${\mathcal{N}}_{p}^{0}$ is selected, meaning the particle only affects the vertices in the cell where it currently resides, and at least ${\mathcal{C}}^{0}$ continuity is required for the kernel. This scenario leads to the kernel degenerating, which is characterized by the kernel affecting merely the shared edge in 2D or the shared face in 3D when the particle transitions across cell boundaries, as depicted in Figure 2a. Conversely, opting for 1-ring neighbors, ${\mathcal{N}}_{p}^{1}$, effectively circumvents this issue, ensuring a non-degenerate kernel interaction as illustrated in Figure 2b.
@@
For conciseness and consistency, we will omit the function arguments and subscripts, such as ${\bm{x}}$ and $i$ in (21). Instead, since we project onto the mesh vertex $v$, we will hereafter use the subscript $v$ and obtain the following rewritten form of (21):
$\begin{bmatrix}\hat{u}_{p}\\ \nabla\hat{u}_{p}\\ \end{bmatrix}={\bm{M}}_{vp}^{-1}{\bm{Q}}_{vp}^{T}{\bm{D}}_{vp}{\bm{u}}_{vp},$ (22)
where ${\bm{u}}_{vp}=\left[u_{v_{1}},\dots,u_{v_{n}}\right]^{T}$ is the stacked vertex values with $v_{i}\in{\mathcal{N}}_{p}^{1}$, $i\in[1,~{}n]$ and the matrix ${\bm{M}}_{vp}={\bm{Q}}_{vp}^{T}{\bm{D}}_{vp}{\bm{Q}}_{vp}$ with stacked vertex basis ${\bm{Q}}_{vp}=\left[{\bm{P}}({\bm{x}}_{v_{1}}-{\bm{x}}_{p}),\dots,{\bm{P}}({\bm{x}}_{v_{n}}-{\bm{x}}_{p})\right]^{T}$ and diagonal sample weighting matrix ${\bm{D}}_{vp}$ such that ${\bm{D}}_{pv,i,i}=d_{vp}=d({\bm{x}}_{v_{1}}-{\bm{x}}_{p})$, where $d$ is chosen to be a B-spline function.
@@ 2.3.3 Required Properties for the Transfer Kernel
Despite implementing a broader interpolation kernel, the reconstruction method described in Section 2.3.2 is not directly applicable as it still suffers from cell crossing errors. To grasp this problem, consider the essential desirable properties for an MPM kernel:
@@
    The kernel must be a non-negative partition of unity. This means that the sum of the kernel weights for all nearby vertices to a particle should equal 1; i.e., $\sum_{v\in{\mathcal{N}}_{p}^{1}}w_{v}=1$, with each individual weight $w_{v}\geq 0,\forall v\in{\mathcal{N}}_{p}^{1}$.
@@
    There should be a continuous reconstruction of both the function value and gradient as the particle transitions across the cell boundary.
Methods such as B-spline MPM and GIMP are specifically designed to fulfill these requirements. With MLS-MPM, the partition of unity is inherently assured by the characteristics of MLS [47]. The non-negativity of this partition additionally depends on preventing sample degeneration, a requirement met in MLS-MPM due to its use of uniform background nodes. Furthermore, MLS-MPM ensures continuous reconstruction by utilizing a B-spline for sample weighting, which provides ${\mathcal{C}}^{1}$ continuity.
A key insight from MLS-MPM is that whenever a grid node is added or removed from the nearby node set, its B-spline weighting function also smoothly approaches zero, ensuring its influence on the assembly of ${\bm{M}}$ and ${\bm{Q}}$ in (21) is infinitesimal and, hence, the continuity of the reconstruction. However, since our method determines nearby vertices based on the ring of neighbors rather than proximity, a uniform sample weighting function cannot guarantee diminishing influence for the added or removed vertices. Consequently, the abrupt changes in influence of these vertices during the MLS assembly yield discontinuous reconstruction. This issue will be addressed in the next section.
@@ 2.3.4 Remedying Discontinuous Reconstruction Across the Cell Boundary
To mitigate the abrupt influence changes from vertices being added or removed during cell crossings, an intuitive solution is to diminish their impact on the MLS assembly. This can be accomplished by adjusting the kernel to predominantly rely on vertices that remain common before and after crossing the cell boundary. To achieve this, we modify any initial sample weighting function $d_{vp}$ by incorporating a smooth diminishing function $\eta_{vp}$; i.e., $d_{vp}\leftarrow\eta_{vp}d_{vp}$. Here, $\eta_{vp}$ approaches zero for vertices that are added or removed from ${\mathcal{N}}_{p}^{1}$ during the cell crossing. A detailed proof of the efficacy of this approach is provided in A.1.
For simplex elements, we design the following $\eta_{vp}$:
$\eta_{vp}=\sum_{n\in{\mathcal{N}}_{p}^{0}}B_{np}{\bm{A}}_{v,n},$ (23)
where, for particle $p$, $B_{np}$ represents the barycentric weight for $n\in{\mathcal{N}}_{p}^{0}$, and ${\bm{A}}$ is the mesh’s adjacency matrix. Geometrically, for a nearby vertex $v\in{\mathcal{N}}_{p}^{1}$, $\eta_{vp}$ denotes the sum of the barycentric weights for all vertices $n\in{\mathcal{N}}_{p}^{0}$ that are directly connected to $v$. Note that $\eta_{vp}=1,~{}\forall v\in{\mathcal{N}}_{p}^{0}$. A.1 proves the claimed diminishing property for this design in the simplex cell.
@@ 2.3.5 Verification of the Proposed Kernel
To verify that the proposed method can produce continuous reconstruction, analytical and numerical solutions of the kernels are produced on 1D and 2D meshes, respectively.
@@
For the 1D case, the first basic verification is conducted on a uniform mesh, as the setup in Figure 3a. Figure 4a shows the correct kernel reconstruction with the diminishing function $\eta$, while Figure 4b, as an ablation, shows that the reconstruction can downgrade to being discontinuous even for the simplest uniform mesh, proving the necessity of $\eta$. The detailed setup for this analytical solution is provided in B.
@@
Note that the presence of a particle within a boundary element, as depicted in Figure 3b, can lead to a negative weight value for the most interior node. This phenomenon is exemplified by Node 3 in Figure 5, where the kernel degenerates due to the absence of a first ring of neighboring elements on the boundary side during MLS sampling. To remedy the problem of negative kernel values, which can potentially lead to numerical instabilities [48], the introduction of an additional layer of elements beyond the original boundary is recommended as a practical solution.
@@
The next verification is on a periodically shrinking and expanding 1D mesh (Figure 3c). The mesh contains cyclic cell sizes of $[\dots,1,R,R^{2},R,1,\dots]$ designed to mimic the transition between varying resolutions. The size transition ratios tested range from $1.1$ to $1.5$ so as to correspond with the typical transition ratios in FEM analysis. Kernel reconstructions are conducted on Nodes 5, 6, 7, and 8 as they can present a full cycle. As shown in Figure 6, both the kernel and the gradient estimations are piece-wise ${\mathcal{C}}^{1}$.
The final verification and ablation tests were performed on a 2D unstructured mesh featuring “&” shapes. The comparison between scenarios with and without the use of $\eta$, as shown in Figure 7a and Figure 7b respectively, validates the importance of $\eta$ and the effectiveness of the proposed method in managing unstructured meshes.
@@ 3 Experiments and Results
To demonstrate and assess the effectiveness of our approach, particularly its reduced cross-cell error owing to the continuous gradient reconstruction, we have chosen representative test cases from prior related studies. Our benchmarking relies on analytical solutions when feasible, or alternatively, on the standard B-spline MPM at a sufficiently high resolution. All experiments were carried out on a single PC equipped with an Intel® Core™ i9-10920X CPU.
@@
[…]
@@
Figure 10a displays the stress profile for a particle located at $x_{0}=12.75$ m, which undergoes the most frequent cell crossings during its vibrational motion. The outcomes achieved with UMLS-MPM showcase a remarkable level of smoothness and precision. Figure 10b illustrates the energy dynamics for the entire system, revealing that the system’s energy is largely conserved throughout the simulation, with only slight fluctuations. These findings collectively underscore the robustness and precision of UMLS-MPM in managing intense cell crossings by particles.
@@
[…]
@@
[…]
@@
[…]
@@
Figure 18: Snapshots of the solutions from the UMLS-MPM and the standard MPM with quadratic B-splines basis functions in Zhao et al. [51]. Particles are colored based on the equivalent plastic strain.
@@
Figure 19: Snapshots of the solutions from the UMLS-MPM and the standard MPM with quadratic B-splines basis functions in Zhao et al. [51]. Particles are colored based on the mean normal stress.
@@
Figure 20: Time evolutions of the run-out distance from UMLS-MPM and the standard MPM.
Figures 18 and 19 show the snapshots of the slope simulated by the standard and UMLS-MPM, where particles are colored by the equivalent plastic strain and mean normal stress, respectively. We can see that UMLS-MPM effectively captures the retrogressive failure pattern of slopes made of sensitive clay. Also, in terms of equivalent plastic strain fields and mean normal stress fields, we observe a strong similarity between the UMLS-MPM solution and the reference solution from [51].

### v2

@@ 2.3.2 Extending MLS-MPM Onto Unstructured Meshes
We select MLS-MPM as our foundation because of its inherent versatility, allowing it to be applied to adjacent nodes without reliance on specific topological or positional constraints. Our implementation and experiments are based on triangular and tetrahedral cells. Nonetheless, it is worth noting that our method can easily be extended to any tessellation by designing a smooth and locally diminishing function $\eta_{v}$ compatible with the tessellation, such as the one in (32) for simplex cells.
@@
To determine the nearby nodes for a given particle $p$, we first locate the cell that encompasses $p$ and refer to its nodes as ${\mathcal{N}}_{p}^{0}$, representing the 0-ring neighbors of $p$. Then, we define ${\mathcal{N}}_{p}^{1}$ as the 1-ring neighbors, which comprise all nodes connected to ${\mathcal{N}}_{p}^{0}$. Note that ${\mathcal{N}}_{p}^{0}\subset{\mathcal{N}}_{p}^{1}$. Similarly, we can define ${\mathcal{N}}_{p}^{2},\dots$ in an analogous manner, as illustrated in Figure 1.
To quickly search for the cell that contains $p$, we pre-store the adjacency relationship between the spatial hashing grid and the mesh cells. When given a ${\bm{x}}_{p}$, the spatial hash grid is queried, and we then only check the cells adjacent to this hash grid. The detailed pipeline can be found in Appendix A.
Figure 2: (a) When ${\mathcal{N}}_{p}^{0}$ alone is selected as the active nearby nodes, as a particle crosses the cell edge, the nodes indicated by the blue and red boxes are added or removed, respectively. Consequently, the weights there must approach zero to ensure ${\mathcal{C}}^{0}$ continuity, resulting in kernel degeneration along the edge. (b) Advancing to ${\mathcal{C}}^{1}$ addresses this issue by incorporating a sufficient number of surrounding nodes to fully encompass the particle.
@@
When a specific level of ring neighbors is chosen as the active set of nodes, a natural question arises:
What is the minimum number of rings required to satisfy the desired properties of the MPM kernel?
Assume ${\mathcal{N}}_{p}^{0}$ is selected, and the particle only affects the nodes $i\in{\mathcal{N}}_{p}^{0}$; since at least ${\mathcal{C}}^{0}$ continuity is required for the kernel, when the particle passes one interface of the cell, the node not on the interface is removed from the active set and the kernel for it must be zero. This leads to the kernel degenerating, i.e., the kernel affecting merely the interface when the particle crosses it, as depicted in Figure 2a. Conversely, opting for 1-ring neighbors, ${\mathcal{N}}_{p}^{1}$, effectively circumvents this issue, ensuring a non-degenerate kernel interaction as illustrated in Figure 2b.
@@
For conciseness, we replace the function input with the subscript, for example replacing $({\bm{x}}_{p})$ with $p$ in all related equations, the reconstruction reads:
$\begin{bmatrix}\hat{u}_{p}\\ \nabla_{{\bm{z}}}\hat{u}_{p}\\ \end{bmatrix}={\bm{M}}^{-1}_{p}{\bm{B}}_{p}{\bm{u}}.$ (31)
@@ 2.3.3 Required Properties for the Transfer Kernel
Consider the essential desirable properties for an MPM kernel:
@@
    The kernel must be a non-negative partition of unity. This means that the sum of the kernel weights for all nearby vertices of a particle should equal 1; i.e., $\sum_{v\in{\mathcal{N}}_{p}^{1}}w_{v}=1$, with each individual weight $w_{v}\geq 0,\forall v\in{\mathcal{N}}_{p}^{1}$.
@@
    There should be a continuous reconstruction of both the function value and gradient as the particle crosses the cell boundary.
With MLS-MPM, the partition of unity is inherently assured by the characteristics of MLS [40], and non-negativity is assured by the uniform sampling of grid nodes (i.e., no degenerate samples). Lastly, with uniform grid nodes, MLS-MPM ensures continuous reconstruction by utilizing a B-spline for sample weighting and provides ${\mathcal{C}}^{1}$ continuity.
However, this property holds only under uniform grid nodes with spacing properly aligned with the support of the B-spline weighting function. The key is that the B-spline function approaches zero for grid nodes that are about to be added or removed. Consequently, the influence of the discrete change in the set of active nodes on the assembly of ${\bm{M}}_{p}$ and ${\bm{B}}_{p}$ in (30) is infinitesimal, ensuring no abrupt change during the reconstruction.
Due to the varying spacing of the unstructured meshes, the weighting function is not guaranteed to approach zero for the added or removed nodes during cell crossings, leading to discontinuous reconstruction. This issue will be addressed in the next section.
Still, we can borrow the key insight from MLS-MPM that “the weighting function for added or removed nodes should approach zero” and design a scheme to enforce this property. The solution will be discussed and presented in the next section.
@@ 2.3.4 Remedying Discontinuous Reconstruction Across the Cell Boundary
The jump change originates from the discrete change of the active set ${\mathcal{N}}_{p}^{1}$ during particle cell crossing if their influence on the MLS assembly is nonzero. Hence, an intuitive solution is to artificially diminish their influence on the MLS assembly. To achieve this, we multiply any initial sample weighting function $d_{p,i}$, such as B-spline, by a smooth diminishing function $\eta_{p,i}$; i.e., $d^{\prime}_{p,i}\leftarrow\eta_{p,i}d_{p,i}$. Here, $\eta_{p,i}\to 0$ for nodes that are added or removed from the active set ${\mathcal{N}}_{p}^{1}$ during the cell crossing. A detailed proof of the efficacy of this approach is provided in Appendix B.
For simplex elements, we design the following $\eta_{p,i}$:
$\eta_{p,i}=\sum_{n\in{\mathcal{N}}_{p}^{0}}B_{p,n}{\bm{A}}_{i,n},$ (32)
where ${\bm{A}}$ denotes the mesh’s adjacency matrix; an adjacency matrix is a binary matrix representing the connectivity of a graph, where each ${\bm{A}}_{i,j}=1$ indicates the presence of an edge between nodes $i$ and $j$. A mesh can naturally be viewed as a graph by connecting an edge between every pair of adjacent nodes in a cell. $B_{p,n}$ represents the barycentric coordinate for particle $p$ with respect to a specific node $n\in{\mathcal{N}}_{p}^{0}$. Taking a 2D simplex, the triangular cell, as an example, the barycentric coordinates of a location ${\bm{x}}_{p}$ are triplets of numbers $b_{1},b_{2},b_{3}$, subject to $b_{1}+b_{2}+b_{3}=1$. If these values are considered as masses placed at the nodes of the triangle, the centroid of these masses will be at ${\bm{x}}_{p}$. Generally, barycentric coordinates can be calculated as follows:
$B_{p,n_{i}}=\frac{\det\left([{\bm{x}}_{n_{1}},\ldots,{\bm{x}}_{n_{i-1}},{\bm{x}}_{p},{\bm{x}}_{n_{i+1}},\ldots,{\bm{x}}_{n_{\text{dim}+1}}]\right)}{\det\left([{\bm{x}}_{n_{1}},\ldots,{\bm{x}}_{n_{\text{dim}+1}}]\right)}.$ (33)
Combining (33) with the adjacency matrix definition provides a more geometric interpretation of the design (32): for $i\in{\mathcal{N}}_{p}^{1}$, $\eta_{p,i}$ is the sum of the barycentric weights for all $n\in{\mathcal{N}}_{p}^{0}$ that are connected to $i$. Note that $\eta_{p,i}=1,~{}\forall v\in{\mathcal{N}}_{p}^{0}$. Appendix B proves the claimed diminishing property for this design in the simplex cell, while Figure 25 provides a simple visual illustration of the proposed $\eta_{p,i}$.
@@ 2.3.5 Verification of the Proposed Kernel
To verify that the proposed method can produce continuous reconstruction, analytical and numerical solutions of some examples are produced in 1D and 2D test cases, respectively.
@@
For the 1D case, the first basic verification is conducted on a uniform mesh, as shown in Figure 3a. Figure 4a shows the correct kernel reconstruction with the diminishing function $\eta$, while Figure 4b, as an ablation, shows that the reconstruction is discontinuous even for the simplest uniform mesh, proving the necessity of $\eta$. The detailed setup for this analytical solution is provided in Appendix C.
@@
Note that when a particle is in a boundary cell, such as Node 3 in Figure 5, negative weight values may be obtained for some interior nodes. This is caused by kernel degeneration due to the absence of a first ring of neighbors on the boundary side during MLS sampling. To remedy this problem, which can cause numerical instabilities [1], an extra layer of cells beyond the real boundary is included in our experiments.
@@
The next verification is on a periodically shrinking and expanding 1D mesh (Figure 3c). The mesh contains cyclic cell sizes of $[\dots,1,R,R^{2},R,1,\dots]$ designed to mimic the transition between varying mesh resolutions. The size transition ratios tested range from $1.1$ to $1.5$ to correspond with typical transition ratios in FEM analysis. Kernel reconstructions are conducted on Nodes 5, 6, 7, and 8 as a full cycle. As shown in Figure 6, both the kernel and the gradient estimations are piece-wise ${\mathcal{C}}^{1}$.
We note that kinks can be seen in the function value plots in Figure 6, leading to the potential confusion that the gradient is discontinuous. However, the plots of kernel values are ${\bm{w}}({\bm{z}};{\bm{x}}_{p})|_{{\bm{z}}={\bm{x}}_{p}}={\bm{w}}({\bm{x}}_{p};{\bm{x}}_{p})$ versus ${\bm{x}}_{p}$, as in (7); for the misleading statement to hold true, the gradient should have been taken with respect to the plot axis ${\bm{x}}_{p}$, i.e., $\nabla_{{\bm{x}}_{p}}{\bm{w}}({\bm{x}}_{p};{\bm{x}}_{p})$, which is not the case according to (8).
The ablation tests are also performed on a 2D unstructured mesh featuring a “&” shape. The comparison between scenarios with and without the use of $\eta$, as shown in Figure 7a and Figure 7b respectively, validates the importance of $\eta$ and the effectiveness of the proposed method in managing unstructured meshes.
Finally, we experimentally show that UMLS-MPM can seamlessly be combined with other schemes, such as the Affine Particle in Cell (APIC) scheme [35, 37] to help conserve the total angular momentum of the system. For details, see Appendix D.
@@ 3 Experiments and Results
To demonstrate and assess the effectiveness of our approach, particularly its reduced cross-cell error owing to the continuous gradient reconstruction, we have chosen representative test cases from prior related studies. Our benchmarking relies on analytical solutions when feasible; alternatively, we use the standard MPM with B-spline or GIMP basis functions at a sufficiently high resolution. All experiments were carried out on a single PC equipped with an Intel® Core™ i9-10920X CPU.
@@
Consider the 1D vibration bar problem shown in Figure 8a [62]. The left end of the bar is fixed and the right has a sliding condition in the $x$ direction. The physical properties of the bar are: $E=100$ Pa, $\nu=0$, $L=25$ m, and $\rho=1$ kg/m3. The initial velocity conditions are $\dot{u}(x,t=0)=v_{0}\sin\left(\beta_{1}x\right)$ with $\beta_{1}=\frac{\pi}{2L}$.
@@
Figure 10a displays the stress profile for a particle located at $x_{0}=12.75$ m, which undergoes the most frequent cell crossings during its vibrational motion. The outcomes achieved with UMLS-MPM showcase a remarkable level of smoothness and precision. Figure 10b illustrates the energy dynamics for the entire system, revealing that the system’s energy is largely conserved throughout the simulation, with only slight fluctuations. We believe the fluctuations in the energy plot are due to the symplectic integration schemes or the combined effect of the FLIP scheme. Similar phenomena have been observed in previous works [16] and [57], respectively. These findings collectively underscore the robustness and precision of UMLS-MPM in managing intense cell crossings by particles.
@@
Next, we considered the problem of two colliding elastic disks shown in Figure 11a [62]. The physical properties of the disks are: $E=1000$ Pa, $\nu=0.3$, $\rho=1000$ kg/m3, and ${\bm{v}}=\pm(0.1,0.1)$ m/s for the left and right disks, respectively. Each disk was discretized with $462$ material points using the triangle mesh of a disk. The background mesh was generated using Delaunay triangulation with a target element size of $0.025$ m. We plot key snapshots of the simulation in Figure 11b–d, with the impact at $1.5$ s, total retardation right before $2.0$ s, and rebounding separation right before $2.5$ s.
@@
Although an unstructured mesh offers the adaptability to match any boundary shape, the cell orientation, or a different tessellation, can potentially affect accuracy. To illustrate the precision of our method under various rotation angles, we examined the case of a cantilever under its own weight, as shown in Figure 13a [62]. The cantilever’s physical characteristics are as follows: length $l=10$ m, height $h=2$ m, gravitational acceleration $g=9.81$ m/s2, Young’s modulus $E=100000$ Pa, Poisson’s ratio $\nu=0.29$, and density $\rho=2$ kg/m3. The cantilever was discretized with uniformly spaced particles in both directions. We created the background mesh using Delaunay triangulation, aiming for an element size of $0.5$ m. Additionally, we rotated the mesh of the cantilever by angles of $15^{\circ}$, $30^{\circ}$, and $45^{\circ}$ to showcase the resilience of our method to rotation, as depicted in Figure 13b–d.
@@
3.4 2D Ball in a Wavy Channel
Figure 17: 2D ball in a wavy channel: snapshots of the simulation at 0s, 1.4s, 3.0s, and 6.5s.
We highlight the proposed method’s ability to conform to irregular geometric boundaries. To this end, we consider the case of a ball freely falling but confined in a wavy channel, as shown in the leftmost subfigure of Figure 17. The physical properties of the ball are: radius $r=1.0$ m, Young’s modulus $E=100$ kPa, Poisson’s ratio $\nu=0.29$, and density $\rho=400$ kg/m3. The ball was discretized with 4735 randomly sampled material points. The background wavy channel has a sinusoidal shape.
The left wall of the channel has an analytical expression of:
$\begin{array}[]{cc}x_{l}=&\begin{cases}A\sin(\omega_{1}y)\sin(\omega_{2}y),&0<y\leq 20.0\text{ m}\\ 0,&\text{Otherwise},\\ \end{cases}\end{array}$ (36)
where $A=1.0$ m, $\omega_{1}=\frac{\pi}{5}$ rad/m, and $\omega_{2}=\frac{\pi}{20}$ rad/m. The right wall is created by shifting the left wall by 2.0 m, i.e., $x_{r}=x_{l}+2.0$ m. The background mesh was generated using Delaunay triangulation with a target element size of 0.05 m, resulting in 43360 cells.
Figure 18: Zoomed-in view of the ball’s deformation in the wavy channel at key timestamps. From top to bottom, timestamps around 1.4s, 3.0s, and 6.0s.
The snapshots of the simulation at 1.4s, 3.0s, and 6.5s are shown in Figure 17, while the zoomed-in views of the ball’s deformation and hydrostatic stress are shown in Figure 18. UMLS-MPM captures both the bouncing into the wavy channel (at 1.1s, 1.9s, 3.7s) and the squeezing through the narrow part of the channel (at 1.4s, 2.4s, 3.0s) with no rasterization artifacts, proving the robustness of the proposed method in simulating under general mesh tessellation and handling irregular geometry boundaries.
@@
Next, the performance of the proposed approach was investigated when dealing with material behavior involving plasticity. To this end, we simulated failure of a 3D slope comprosed of sensitive clay. The problem geometry was adopted from [69] and is illustrated in Figure 19. Here, the bottom boundary of the slope is fixed and the three lateral sides are supported with rollers. To model the elastoplastic behavior of the sensitive clay in an undrained condition, a combination of Hencky elasticity and J2 plasticity with softening was used. The softening behavior is governed by the following exponential form: $\kappa=(\kappa_{p}-\kappa_{r})e^{-\eta\varepsilon_{q}^{p}}+\kappa_{r}$, where $\kappa$, $\kappa_{p}$, and $\kappa_{r}$ denote the yield strength, the peak strength, and the residual strength, respectively, $\varepsilon_{q}^{p}$ denotes the equivalent plastic strain, and $\eta$ is a softening parameter. The specific parameters were adopted from [69]. They are a Young’s modulus of $E=25$ MPa, a Poisson’s ratio of $\nu=0.499$, a peak strength of $\kappa_{p}=40.82$ kPa, a residual strength of $\kappa_{r}=2.45$ kPa, and a softening parameter of $\eta=5$. The assigned soil density is $\rho=2.15$ t/m3.
@@
Figure 20: Snapshots of the solutions from the UMLS-MPM and the standard MPM with GIMP basis functions in Zhao et al. [69]. Particles are colored based on the equivalent plastic strain.
@@
Figure 21: Snapshots of the solutions from the UMLS-MPM and the standard MPM with GIMP basis functions in Zhao et al. [69]. Particles are colored based on the mean normal stress.
@@
Figure 22: Time evolutions of the run-out distance from UMLS-MPM and the standard MPM with GIMP basis functions.
Figures 20 and 21 show the snapshots of the slope simulated by the standard and UMLS-MPM, where particles are colored by the equivalent plastic strain and mean normal stress, respectively. We can see that UMLS-MPM effectively captures the retrogressive failure pattern of slopes made of sensitive clay. Also, in terms of equivalent plastic strain fields and mean normal stress fields, we observe a strong similarity between the UMLS-MPM solution and the reference solution from [69].